What is the mass of 2.51 × 1024 atoms (particles) of lead metal?


1
Expert's answer
2021-03-30T07:56:48-0400

= 2.51 × 10^24/ 6.022 × 10^23

= 0.4168 × 10^1 = 4.618mol Pb

= 4.618 × (207.2 g/mol)

= 956.8496g of lead metal